by TEI TEI

05/03/07 – After we found that the existing expansion tank was ruptured we did some super quick engineering and worked it out with PHB to install a new bladder tank Expansion tank on the 10th floor set to 20 psig. After some initial heartache, we found this fix to be working out nicely.

03/15/07 – Joe reported that pump head was extremely high in low load conditions. TE to investigate.

02/19/07 – PHB has finished the pipe bridge work and corrected the reverse return system at the building. First test appear to have everything working correctly. TEI directed everyone to closeout the project.

02/02/07 – PHB is having Bob Gettler’s brother come down to the building next monday to look over remaining work to take Bob’s place.

01/17/07 – TE distributed the punchlist and informed contractors to turn in all closeout paperwork and final pay applications for this project. Punchlist PDF

01/05/07 – TE has to get Trane and PHB to fix a couple of items on the chiller installation and some prices to finish the work including a valve in the pipe bridge and some steps.

12/04/06 – TTC informed contractors to use 10/11/2006 for warranty start date on this project. It was the weekend of the startup and electrical shutdown and will be as good as any to use as the chiller was brought online for service shortly after that. After talking to Joe, we may want PHB to give us some prices for steps and rails up to the chiller control panel as well as the other RTU’s with our allowance money.

12/01/06 – TTC talked to QCE and they will have O&M’s here in a week. TTC has approved payapp #2 for this project for them and sent to Tony.

10/25/06 – PHB Change Order for piping change will come out of allowance of $5000 which has not been used yet (see below). Queen City also confirmed that the $5000 allowance was not used to date.

10/10/06 – TTC check on weekend shutdown and everything seemed to go well. The chiller was not actually started. Today, Bernie told TEI that he believed QCE would have a person down at the Admin building to actual provide power to the chiller. This will allow the chiller to be tested and factory started on Wednesday as planned. TTC may stop down in the morning to check on progress.

10/02/06 – Per everyone’s agreement the chiller will be started on 10/11/06.

09/01/06 – Greg reported today that Duke has been inside the vault and working on the new service. This could mean a completion time of next week. QCE will follow that up with a scheduled shutdown of the building (after talking to Joe) for their tie-in. PHB will be down next week to finalize the piping bridge crossover removal work and then we can really wrap this project up. The new chiller will need to be powered up and tested and the controls (current status unknown by TTC) will have to be checked. Any new computer room system can be added to the system from that point on as long as everyone understands that chilled water will now be available in the building 24/7 either my the main plant or the rooftop plant.

08/18/06 – TTC emailed Joe and Tony costs on converting the system to 24/7 chilled water for computer room units. It can be done with little up front cost to the building and each user would pay for connection to the system. Each computer room unit would get a tap on its associated floor, a check valve, an isolation valve, and an inline pump to provide the chilled water to the unit.

06/29/06 – TTC met Bob Gettler on site to discuss the removal of the bridge connecting the supply and return piping for the CHWS/R. PHB will put a price and time together to do this as Joe and I decided it should be done on the weekend and some equipment should be moved to keep it from getting wet. They will put a price together and send it to us for approval.

06/08/06 -TTC brought all the equipment to the site for field work. We took many readings, some of which don’t make sense. The pressure drop across the flow meter in the basement is showing a chiller gpm flow that is 100 gpm higher than design with or without the auxiliary pump on. The pressure drop at the top of the riser is about 30 psig but the pressure drop across the coil is 1.5 psig. The water temperature at the inlet to the coil is only 57.4 degrees. Cooling is an issue on this floor. The bridge could have something to do with it. TTC requested data from Trane and the balance report to help in resolving the new temperature issues. TZ remembered that during design, this floor had temperature issues during hot weather too so that may be indicative of a long term problem.

06/06/06 – TTC and Joe have been working on AHU-10 serving the 8th floor space since the flow and temperature don’t seem to be exact. Joe also showed TTC a piping bridge that existing on the existing supply and return piping that wasn’t shown on the TEI PID. After consultation with Todd Zurweller and Gary Davies, it was agreed that the new design cannot work without that bridge being removed and blocked. TTC will stop by on 6-7-06 to field verify flows and pressures. TTC emailed Tony with Duke Energy contact for new electric meter application. QCE is negotiating with PHB on roof curb space for new conduits. TTC emailed PHB about removing the 10th floor bridge.

05/26/06 – The steel beam has been moved to its correct location. TTC gave PHB the okay to remove the expansion tank as part of the allowance as long as it is under $1000 as we have a $5000 allowance on this project for the MEP contractor. This expansion tank will no longer be needed on this project but it was not shown as demolished on the plans – it was left as is. TTC contact QCE to have the remove two conduits that are in the way of finishing the demo of the existing chiller.

0523/06 – The chiller lift went very well aside from the new support steel being 1-1/2″ off from where it should have been welded. THP, PHB, TEI and HamCo met onsite to discuss options. TTC instructed PHB to move the steel to the correct location. THP provided a field sketch showing how to use a spacer to keep extra work down to a minimum. PDF. The old chillers were removed and the new chiller set in place with little other incident. PHB pointed out about the expansion tank in the 10th floor. It is not necessary, Joe would like it removed, but the TE dwgs do not show it being demolished. TTC asked PHB for an additional cost for this work. PHB pointed out two conduits they will need removed and TTC said he would get that taken care of by QCE. The electric penetrations for this job were also discussed and TTC said that QCE could use the refrigerant penetrations to save additional roof penetrations and TTC will talk to QCE about this. Lastly, PHB asked about routing the new chilled water lines out of the mechanical room in a different spot closer to the steel. TTC approved this new location as he did not see negatives to the proposed locations.

05/18/06 – PHB and TEI are discussing Saturday Chiller lift. PHB asked about vibration isolators. They do not come with the chiller but TTC would like them to be put under the chiller before it is set on the ground. We asked for a cost. It should only be a couple of hundred bucks.

05/09/06 – Const Mtg 2 was held and mtg notes distributed. The chiller list is scheduled for May 20.

03/29/06 – TTC distributed meeting minutes from yesterdays meeting. We are shooting for do the weekend shutdown over the May 20 weekend if the chiller arrives the week of May 15 like it is supposed too.

03/20/06 – City asked for stamped structural dwgs on this project. TTC had THP stamp their drawings and send it over for PHB to replace in the permit set at the city.

03/15/06 – TEI approved chiller, pump, wiring, mat and switchgear submittals and sent them out to the contractors and owner3/09/06 – Chiller and Electrical submittals have arrived. The chiller selection will be modified for 25% glycol. This does not adversely affect the chiller selection by more than 5 tons. PHB however will need to add some glycol to the system when it is refilled. Also pressure taps on the top of the building should be routed to a drum for collection. These are extra cost to the project that my have to be incorporated. TZ and TTC decided that keeping the electric heat tracing on the piping is still a good idea since it is already included in the project.

02/27/06 – Pre Construction minutes sent to contractors. PHB submittals arrived today. TEI did not select this chiller on glycol. TEI has to contact the manufacturer about derating the chiller for glycol and see what the final tonnage is for this building.

02/09/06 – Pre Construction meeting held.

12/13/05 – Contracts received, corrected and shipped back to the County.

12/01/05 – Trane called to tell everyone that prices were increasing this day and they want a confirmed order. TTC and Tony told Bernie at PHB that they could hold the unit for us while we wait for final contract approval. It was noted that the bid is still valid for 60 days so this price increase should not affect our project. Trane should be made aware.

11/04/05 – TTC, Tony and Bernie finalized bid numbers and some issues with was actually submitted to the County on bid day. Tony gave TE the okay to write the contract for this project. TE will have contract finished by Wed of next week.

10/24/05 – Final Addendum went out last Friday for project. Bid opening is set for this Thursday.

10/14/05 – TTC left a requirement about testing in the spec that has to be removed. Single compressors on each circuit are the only one that York makes so TTC will have to pick that up in the Addendum. The prebid meeting was successful.

10/04/05 – TE has finished the drawings and specs for the rebid and is sending it out for bidding to HamCo Purchasing.

10/03/05 – IBI confirmed today that with the 800A breaker on the 10th floor we can buy the 150-ton chiller as long as the MCA is less than 800A. Our 150 ton chillers are 733A so we can go with the full sized chiller. TTC and MS will finish up the drawings today and will deliver them to the County on Tuesday.

09/27/05 – Tony gave TE the okay to go with a 125-ton chiller if needed. In talking with George Mullins and Mike Stiles today, there may still be a change to stay with the 150-ton chiller because the total cost of the electrical project has decreased from a $105,000 job to a $25,000 project. With this additional money we can upgrade the GE chiller disconnect to 1200A. This will allow us to increase the chiller size back to 150-ton. Mike and George have to setup a field trip with Joe Merkt to verify some electrical data to make this happen. We are still trying to get the rebid done by this Friday.

09/18/05 – ThermalTech began re-engineering the project for buying the chillers on LCC and re-doing the electric design to meet the new standard for providing dual power to all equipment in the building. We want to be complete with the rebid this Friday but are running into some roadblocks where it concerns feeding the new rooftop chiller with an 800A max service or MOCP.